# MGMCpaper.bib

### v2

@@ Abstract
Gaussian random fields play an important role in many areas of science and engineering. In practice, they are often simulated by sampling from a high-dimensional multivariate normal distribution, which arises from the discretisation of a suitable precision operator. Existing methods such as Cholesky factorization and Gibbs sampling become prohibitively expensive on fine meshes due to their high computational cost. In this work, we revisit the Multigrid Monte Carlo (MGMC) algorithm developed by Goodman & Sokal (Physical Review D 40.6, 1989) in the quantum physics context. To show that MGMC can overcome these issues, we establish a grid-size-independent convergence theory based on the link between linear solvers and samplers for multivariate normal distributions, drawing on standard multigrid convergence theory. We then apply this theory to linear Bayesian inverse problems. This application is achieved by extending the standard multigrid theory to operators with a low-rank perturbation. Moreover, we develop a novel bespoke random smoother which takes care of the low-rank updates that arise in constructing posterior moments. In particular, we prove that Multigrid Monte Carlo is algorithmically optimal in the limit of the grid-size going to zero. Numerical results support our theory, demonstrating that Multigrid Monte Carlo can be significantly more efficient than alternative methods when applied in a Bayesian setting.
@@ 1 Introduction
The approximate simulation of Gaussian random fields plays a pivotal role in a large number of research areas, such as quantum physics [goodman1989multigrid, edwards1992multi], spatial statistics [LindgrenEtAl.F_2011_ExplicitLinkGaussian], additive manufacturing [ZhangEtAl.H_2021_StochasticModelingGeometrical], cosmology [Marinucci.D_2011_book], natural language processing [inoue-etal-2022-infinite], public health [MannsethEtAl.J_2021_VariationUseCaesarean], geosciences [LiuEtAl.Y_2019_AdvancesGaussianRandom], or uncertainty quantification in engineering applications [Soize.C_2017_UncertaintyQuantificationAccelerated, VasileEtAl.M_2021_AdvancesUncertaintyQuantification]. In theory, it is a well understood and solved problem, but in practice the efficiency and performance of traditional sampling procedures degenerates quickly when the random field is discretised on a grid (lattice) with spatial resolution going to zero. Moreover, most existing algorithms do not scale well on large-scale parallel computers. Direct approaches based on factorisations of the covariance or precision matrix, e.g., [rue2001fast, rue2005gaussian], run into memory problems and their cost per sample scales in general like $\mathcal{O}(n^{1+\zeta})$ with respect to the number of grid points $n$ where $\zeta>0$ is a positive constant; the computational complexity can only be reduced to $\mathcal{O}(n\log(n))$ if the matrix is structured and FFT-based approaches are applicable. Additional complications arise in the case where the Gaussian distribution is conditioned on data, as in (linear) Bayesian inverse problems, or when the spectrum of the covariance operator decays slowly. On the other hand, stationary, iterative approaches, such as random-walk Metropolis-Hastings [metropolis1953, hastings1970], preconditoned Crank–Nicholson [cotter2013mcmc], Langevin-based samplers [roberts1996exponential] or Hamiltonian Monte Carlo [duane1987216] become extremely inefficient as the grid size goes to zero. In particular, the convergence rate and the integrated autocorrelation time of the resulting Markov chain degenerate, leading to poor effective sample sizes. A lot of work has gone into this problem, but it is still a topic of ongoing research [LindgrenEtAl.F_2011_ExplicitLinkGaussian, LiuEtAl.Y_2019_AdvancesGaussianRandom, LangEtAl.A_2011_FastSimulationGaussian, Croci.M_etal_2018_EfficientWhiteNoise, FeischlEtAl.M_2018_FastRandomField, BachmayrEtAl.M_2020_UnifiedAnalysisPeriodizationBased, BolinEtAl.D_2020_RationalSPDEApproach, KressnerEtAl.D_2020_CertifiedFastComputations, LindgrenEtAl.F_2022_SPDEApproachGaussian].
@@
One of the most basic stationary iterative methods is the Gibbs sampler, see e.g. [RobertEtAl.CP_2004_MonteCarloStatistical, GelmanEtAl.A_2013_BayesianDataAnalysis, RubinsteinEtAl.RY_2016_SimulationMonteCarlo]. The problem of this standard method is its slow convergence to the target distribution and the strong correlation of the resulting samples, both of which become substantially worse as the resolution increases. Already in the late 1980s, Goodman, Sokal and their collaborators wrote a series of papers aimed at accelerating the Gibbs sampler using multigrid ideas, leading to extensive research activity in the quantum physics community [goodman1989multigrid, edwards1992multi, edwards1991multi, mana1996dynamic, mendes1996multi]. The key observation in the seminal paper by Goodman and Sokal [goodman1989multigrid] is the connection of random samplers, such as the Gibbs method, to iterative solvers for linear systems. In particular, the Gibbs sampler for generating Gaussian samples $x\sim\mathcal{N}(\overline{x},A^{-1})$ with mean $\overline{x}$ and covariance matrix $A^{-1}$ can be written as an iteration identical, except for an additive noise component, to the Gauss–Seidel iteration for solving the linear system $A\overline{x}=f$ to obtain $\overline{x}=A^{-1}f$; for more details see [goodman1989multigrid], as well as [Fox.C_Parker_2017_AcceleratedGibbsSampling]. Based on this observation, they proposed the so-called Multigrid Monte Carlo (MGMC) method—a random analogue of the multigrid method for solving discretised partial differential equations (PDEs). A careful analysis of the MGMC method is the focus of this paper. We assume that we have a positive definite operator $\mathcal{A}:H\to H$ on some function space $H$ on $\mathbb{R}^{d}$, for example $\mathcal{A}=(-\Delta+\kappa^{2}I)^{\alpha}$ with positive integer $\alpha$ where $\Delta$ is the Laplace operator. A sufficiently accurate discretisation of $\mathcal{A}$ on some grid in $\mathbb{R}^{d}$ results in a possibly very large matrix $A\in\mathbb{R}^{n\times n}$. The goal of this paper is to show how to efficiently generate samples from $\mathcal{N}(\overline{x},A^{-1})$ for $n\rightarrow\infty$. We provide a rigorous theoretical justification for the optimal scaling of our method for large scale problems: the cost for generating an independent sample grows linearly with the problem size and the sampler is optimal in the continuum limit. Our analysis includes the important setting of a Gaussian random field conditioned on noisy data via a Bayesian approach; here, $\mathcal{A}$ is a finite-rank perturbation of some differential operator and the quantity of interest a functional of the Gaussian random field. Somewhat surprisingly, MGMC has not yet been considered in this context.
@@
    A grid-size-independent convergence theory for the MGMC method. The analysis shows that the first two moments, which fully characterise the Gaussian distribution, converge exponentially at a uniform rate to the target moments.
@@
    A corollary on the optimal complexity of MGMC to produce approximate samples whose mean and covariance converge to the mean and covariance of the limiting infinite-dimensional Gaussian random field.
@@
In contrast to stationary iterative methods on a single grid, such as the Gibbs sampler, the convergence rate of MGMC does not degenerate as the grid size tends to zero. The convergence analysis follows Fox and Parker [Fox.C_Parker_2017_AcceleratedGibbsSampling]. In that paper, they demonstrate that the convergence of the standard component-sweep Gibbs sampler is (in a suitable sense) equivalent to the deterministic (lexicographical) Gauss–Seidel iterative method, and that any matrix splitting used to generate a deterministic relaxation scheme also induces a stochastic relaxation that is a generalised Gibbs sampler; see also [norton2016tuning] where this analysis is extended also to non-Gaussian targets and to the popular Langevin and hybrid (or Hamiltonian) Monte Carlo samplers. Here, we extend the analysis in [Fox.C_Parker_2017_AcceleratedGibbsSampling] to MGMC. In doing so, we need to go beyond the derivation in [goodman1989multigrid] and beyond the standard multigrid theory (as presented for example in [Hackbusch.W_2016_IterativeSolutionLarge]) which provides convergence of the mean. Extra work is necessary to obtain the convergence of the covariance and to extend the analysis to the conditional case in a Bayesian setting. In particular, for linear functionals of the random field that are bounded in $H$, the crucial approximation and smoothing properties for the conditional case can be reduced to the corresponding properties in the unconditional case.
@@
The link between stationary iterative methods from numercial linear algebra and generalised Gibbs samplers was first noted and exploited by Adler and Whitmer [adler1981over, whitmer1984over] and then by Goodman and Sokal [goodman1989multigrid] in the context of MGMC, before being formalised in a mathematically abstract way and applied to linear Bayesian inverse problems in [Fox.C_Parker_2017_AcceleratedGibbsSampling]. A key observation in [Fox.C_Parker_2017_AcceleratedGibbsSampling] is that stationary iterative methods, such as Gauss-Seidel, might have been investigated in the 1950s as linear solvers, but their convergence is too slow to be of any practial interest for solving discretised PDEs in current days. The often purported slow convergence of MCMC methods based on stationary iterative procedures such as the Gibbs sampler thus follows directly from said equivalence between stationary iterative solvers and generalised Gibbs samplers. Parker and Fox then go on to note that “the last fifty years has seen an explosion of theoretical results and algorithmic development that have made linear solvers faster and more efficient, so that for large problems, stationary methods are used as preconditioners at best, while the method of preconditioned conjugate gradients, GMRES, multigrid, or fast-multipole methods are the current state-of-the-art for solving linear systems in a finite number of steps [saad2000iterative]. In [Fox.C_Parker_2017_AcceleratedGibbsSampling] and the earlier papers [parker2012sampling, fox2014convergence], they exploit this link and demonstrate how to substantially speed up the stationary iterations derived from symmetric splittings by polynomial acceleration, particularly Chebyshev acceleration or Krylov-type methods such as conjugate gradients. However, the authors of [Fox.C_Parker_2017_AcceleratedGibbsSampling] do not go on to analyse the MGMC method by Goodman and Sokal or to extend it to the linear Bayesian setting that we consider here.
@@
A range of other methods have been proposed for sampling from multivariate normal distributions that arise from the discretisation of Gaussian random fields; e.g. [rue2001fast, rue2005gaussian] describe several optimisations of the Cholesky sampler for problems that are formulated on a graph. A suitable reordering of the precision matrix results in a reduced bandwidth and thus a much better computational complexity. As the authors show, the method can be further optimised by recursively sampling from conditional distributions that arise from partitioning the graph via a divide-and-conquer approach; this requires the solution of sparse linear systems for which standard methods can be employed.
@@
The precision operator $\mathcal{A}$ and the dimension $d$ of the problem determine the roughness of the underlying Gaussian Random Field. The commonly used operator $\mathcal{A}=(-\Delta+\kappa^{2}I)^{\alpha}$ corresponds to a class of Matern fields with $\nu=\alpha-d/2$ where $\nu$ determines the mean-square differentiability of Gaussian Random Field [LindgrenEtAl.F_2011_ExplicitLinkGaussian]. Although the discretised problem is well-defined for a finite lattice spacing, for $\nu\leq 0$ the field is so rough that the continuum limit of the probability distributions only exists in the weak sense, i.e. when evaluated against bounded linear operators. This includes the important case $\mathcal{A}=-\Delta+\kappa^{2}I$ in $d>1$ dimensions. In contrast to other methods such as the SPDE approach [LindgrenEtAl.F_2011_ExplicitLinkGaussian, LindgrenEtAl.F_2022_SPDEApproachGaussian, Croci.M_etal_2018_EfficientWhiteNoise], this does not pose any problems for MGMC. To see this, observe that both MGMC and the SPDE approach work with the precision matrix $A$. This leads to good computational properties when $A$ is sparse, and this naturally follows if the matrix is given by a discretised differential operator of the above form. However, the SPDE approach inverts the square-root $A^{1/2}$ of the precision matrix, which corresponds to the operator $\mathcal{A}^{1/2}=(-\Delta+\kappa^{2}I)^{\alpha/2}$. If $\alpha$ is an odd integer, the SPDE approach will require the inversion of a fractional power of a sparse matrix, which is not sparse itself. This problem does not arise for MGMC which only uses $A$ and which can therefore readily be used to sample fields for $\mathcal{A}=-\Delta+\kappa^{2}I$. More generally, for arbitrary precision operators MGMC is therefore able to also cover the case where the square-root of the precision matrix is non-sparse even though the precision matrix itself is sparse.
@@
In [harbrecht2012low] a low-rank truncation of the Cholesky factorisation is used to approximately sample from Gaussian random fields. Although the method introduces an approximation, the resulting error can be controlled systematically and the method can be more efficient than the naive Cholesky approach since only a finite number of eigenmodes have to be included. However, the efficiency of the method depends on the spectrum of the covariance operator, and in particular the sufficiently rapid decay of the eigenvalues. We have found that for $\mathcal{A}=-\Delta+\kappa^{2}I$ in $d>1$ dimensions the number of required eigenmodes grows too rapidly for the method to be efficient. Again, MGMC is able to cover this case without any problems.
@@
Another advantage of the MGMC approach is that it can be readily parallelised on distributed memory machines by using well established domain decomposition techniques for the parallel implementation of multigrid solvers [chow2006survey]. For example, the grid traversal in the Gibbs sampler that is used at every level of the grid hierarchy, needs to be performed in a red-black ordering and well-known strategies such as coarse grid aggregation [blatt2012massively] can be used on the coarser levels to take into account the fact that the number of processing units can exceed the number of unknowns. Sampling from posterior distributions in the linear Bayesian setting that we consider here, requires parallel sparse matrix vector products and scatter operations of the form $y=a(a^{\top}x)$ for vectors $a,x,y$ and again these can be implemented using well established techniques in scientific computing. Although a parallel implementation of MGMC is beyond the scope of this paper, we have every reason to believe that it will show the same excellent parallel scalability on large distributed memory machines as multigrid solvers [baker2012scaling, blatt2012massively, kohl2022textbook].

### v3

# Multigrid Monte Carlo Revisited: Theory and Bayesian Inference

@@ Abstract
Gaussian random fields play an important role in many areas of science and engineering. In practice, they are often simulated by sampling from a high-dimensional multivariate normal distribution, which arises from the discretisation of a suitable precision operator. Existing methods such as Cholesky factorization and Gibbs sampling become prohibitively expensive on fine meshes due to their high computational cost. In this work, we revisit the Multigrid Monte Carlo (MGMC) algorithm developed by Goodman & Sokal (Physical Review D 40.6, 1989) in the quantum physics context. While the authors of this paper conclude that MGMC does not overcome critical slowing down in simulations of field theories near phase transitions, we demonstrate here that it has the potential to significantly accelerate sampling in spatial statistics. The class of Gaussian Random Fields we consider includes those with Matérn covariance, but is more general in that it also allows for non-stationary covariance functions. To show that MGMC can overcome the limitation of existing methods, we establish a grid-size-independent convergence theory based on the link between linear solvers and samplers for multivariate normal distributions, drawing on standard multigrid convergence arguments. We then apply this theory to linear Bayesian inverse problems. This application is achieved by extending the standard multigrid theory to operators with a low-rank perturbation. Moreover, we develop a novel bespoke random smoother which takes care of the low-rank updates that arise in constructing posterior moments. In particular, we prove that Multigrid Monte Carlo is algorithmically optimal in the limit of the grid-size going to zero. Numerical results support our theory, demonstrating that Multigrid Monte Carlo can be significantly more efficient than alternative methods when applied in a Bayesian setting.
@@ 1 Introduction
The approximate simulation of Gaussian random fields plays a pivotal role in a large number of research areas, such as quantum physics [undef, undefa], spatial statistics [undefb], additive manufacturing [undefc], cosmology [undefd], natural language processing [undefe], public health [undeff], geosciences [undefg], or uncertainty quantification in engineering applications [undefh, undefi]. In theory, it is a well understood and solved problem, but in practice the efficiency and performance of traditional sampling procedures degenerates quickly when the random field is discretised on a grid (lattice) with spatial resolution going to zero. Moreover, most existing algorithms do not scale well on large-scale parallel computers. Direct approaches based on factorisations of the covariance or precision matrix, e.g., [undefj, undefk], run into memory problems and their cost per sample scales in general like $\mathcal{O}(n^{1+\zeta})$ with respect to the number of grid points $n$ where $\zeta>0$ is a positive constant; the computational complexity can only be reduced to $\mathcal{O}(n\log(n))$ if the matrix is structured and FFT-based approaches are applicable. While in exact arithmetic the Cholesky factorisation allows the generation of independent samples from the target distribution, this is not the case in finite precision floating point arithmetic: for very large problems the ill-conditioning of the precision matrix can lead to the accumulation of rounding errors and the generated samples might in fact be drawn from a slightly biased distribution.
Additional complications arise in the case where the Gaussian distribution is conditioned on data, as in (linear) Bayesian inverse problems, or when the spectrum of the covariance operator decays slowly. On the other hand, stationary, iterative approaches, such as random-walk Metropolis-Hastings [undefl, undefm], preconditoned Crank–Nicholson [undefn], Langevin-based samplers [undefo] or Hamiltonian Monte Carlo [undefp] become extremely inefficient as the grid size goes to zero. In particular, the convergence rate and the integrated autocorrelation time of the resulting Markov chain degenerate, leading to poor effective sample sizes. A lot of work has gone into this problem, but it is still a topic of ongoing research [undefb, undefg, undefq, undefr, undefs, undeft, undefu, undefv, undefw].
@@
One of the most basic stationary iterative methods is the Gibbs sampler, see e.g. [undefx, undefy, undefz]. The problem of this standard method is its slow convergence to the target distribution and the strong correlation of the resulting samples, both of which become substantially worse as the resolution increases. Already in the late 1980s, Goodman, Sokal and their collaborators wrote a series of papers aimed at accelerating the Gibbs sampler using multigrid ideas, leading to extensive research activity in the quantum physics community [undef, undefa, undefaa, undefab, undefac]. The key observation in the seminal paper by Goodman and Sokal [undef] is the connection of random samplers, such as the Gibbs method, to iterative solvers for linear systems. In particular, the Gibbs sampler for generating Gaussian samples $x\sim\mathcal{N}(\overline{x},A^{-1})$ with mean $\overline{x}$ and covariance matrix $A^{-1}$ can be written as an iteration identical, except for an additive noise component, to the Gauss–Seidel iteration for solving the linear system $A\overline{x}=f$ to obtain $\overline{x}=A^{-1}f$; for more details see [undef], as well as [undefad, undefae]. Based on this observation, they proposed the so-called Multigrid Monte Carlo (MGMC) method—a random analogue of the multigrid method for solving discretised partial differential equations (PDEs). A careful analysis of the MGMC method is the focus of this paper. We assume that we have a positive definite operator $\mathcal{A}:H\to H$ on some function space $H$ on $\mathbb{R}^{d}$, for example $\mathcal{A}=(-\Delta+\kappa^{2}I)^{\alpha}$ with positive integer $\alpha$ where $\Delta$ is the Laplace operator. A sufficiently accurate discretisation of $\mathcal{A}$ on some grid in $\mathbb{R}^{d}$ results in a possibly very large matrix $A\in\mathbb{R}^{n\times n}$. The goal of this paper is to show how to efficiently generate samples from $\mathcal{N}(\overline{x},A^{-1})$ for $n\rightarrow\infty$. It should be pointed out that a discussion of the (rate of) convergence of the finite dimensional distribution $\mathcal{N}(\overline{x},A^{-1})$ to the corresponding inifinite dimensional Gaussian Random Field is beyond the scope of this paper, and we refer the reader to [undefw] and references therein for further details. We provide a rigorous theoretical justification for the optimal scaling of our method for large scale problems: the cost for generating an independent sample grows linearly with the problem size and the sampler is optimal in the continuum limit. Our analysis includes the important setting of a Gaussian random field conditioned on noisy data via a Bayesian approach; here, $\mathcal{A}$ is a finite-rank perturbation of some differential operator and the quantity of interest a functional of the Gaussian random field. Somewhat surprisingly, MGMC has not yet been considered in this context.
@@
While the focus of this paper is the application of MGMC in a geometric setting which naturally induces a hierarchy of nested function spaces, the method is more general and can also be applied in a purely algebraic sense. This extension to an algebraic multigrid Monte Carlo method is beyond the scope of this paper and the topic of a subsequent publication, based on preliminary work in [undefaf].
@@
    A grid-size-independent convergence theory for the MGMC sampler. The analysis shows that the first two moments of the samples, which fully characterise the Gaussian distribution, converge exponentially at a uniform rate to the target moments.
@@
    A corollary on the root mean squared convergence of a Monte Carlo estimator that uses the samples generated by the MGMC algorithm to predict some quantity of interest.
@@
    A corollary on the optimal complexity of MGMC updates to produce samples, whose mean and covariance converge to those of the limiting infinite-dimensional Gaussian random field.
  6. 6.
@@
In contrast to stationary iterative methods on a single grid, such as the Gibbs sampler, the convergence rate of MGMC does not degenerate as the grid size tends to zero. The convergence analysis follows Fox and Parker [undefae]. In that paper, they demonstrate that the convergence of the standard component-sweep Gibbs sampler is (in a suitable sense) equivalent to the deterministic (lexicographical) Gauss–Seidel iterative method, and that any matrix splitting used to generate a deterministic relaxation scheme also induces a stochastic relaxation that is a generalised Gibbs sampler; see also [undefag] where this analysis is extended also to non-Gaussian targets and to the popular Langevin and hybrid (or Hamiltonian) Monte Carlo samplers. Here, we extend the analysis in [undefae] to MGMC. In doing so, we need to go beyond the derivation in [undef] and beyond the standard multigrid theory (as presented for example in [undefah]) which provides convergence of the mean. Extra work is necessary to obtain the convergence of the covariance and to extend the analysis to the conditional case in a Bayesian setting. In particular, for linear functionals of the random field that are bounded in $H$, the crucial approximation and smoothing properties for the conditional case can be reduced to the corresponding properties in the unconditional case.
@@
The link between stationary iterative methods from numercial linear algebra and generalised Gibbs samplers was first noted and exploited by Adler and Whitmer [undefai, undefaj] and then by Goodman and Sokal [undef] in the context of MGMC. In [undefak] the convergence of random sampling schemes [^1] In [undefak], a “deterministic” update rule sweeps through the state vector in a fixed order and updates local values by drawing from a random distribution. In contrast, “stochastic” updates process the coordinates in random order. is analysed and bounds for the rate of convergence are derived for Gaussian probability distributions. The authors of [undefak] point out that the key quantity in the analysis is the error-propagation matrix of the Gauss-Seidel iteration. The work in [undefal, undefam] further extends the analysis presented in [undefak], while [undefan, undefao] discuss stochastic relaxation methods and the Gibbs sampler in the context of image processing. However, as far as we can tell these papers do not explicitly point out the connection between Gibbs samplers and relaxation methods for the solution of linear systems. As [undefan] remarks, a random sampler that only allows transitions to states with lower energy will get stuck in local minima. Of course, for a multivariate normal distribution this implies convergence to the most likely state, which is the solution of a linear system.
@@
The link between iterative methods and samplers was formalised in a mathematically abstract way in [undefae]. They generalised the convergence analysis to any sampler that corresponds to a general matrix splitting and applied the accelerated sampler to linear Bayesian inverse problems. A key observation in [undefae] is that stationary iterative methods, such as Gauss-Seidel, might have been investigated in the 1950s as linear solvers, but their convergence is too slow to be of any practial interest for solving discretised PDEs in current days. The often purported slow convergence of MCMC methods based on stationary iterative procedures such as the Gibbs sampler thus follows directly from said equivalence between stationary iterative solvers and generalised Gibbs samplers. Parker and Fox then go on to note that “the last fifty years has seen an explosion of theoretical results and algorithmic development that have made linear solvers faster and more efficient, so that for large problems, stationary methods are used as preconditioners at best, while the method of preconditioned conjugate gradients, GMRES, multigrid, or fast-multipole methods are the current state-of-the-art for solving linear systems in a finite number of steps [undefap]. In [undefae] and the earlier papers [undefaq, undefar], they exploit this link and demonstrate how to substantially speed up the stationary iterations derived from symmetric splittings by polynomial acceleration, particularly Chebyshev acceleration or Krylov-type methods such as conjugate gradients. However, the authors of [undefae] do not go on to analyse the MGMC method by Goodman and Sokal or to extend it to the linear Bayesian setting that we consider here.
@@
A range of other methods have been proposed for sampling from multivariate normal distributions that arise from the discretisation of Gaussian random fields; e.g. [undefj, undefk] describe several optimisations of the Cholesky sampler for problems that are formulated on a graph. A suitable reordering of the precision matrix results in a reduced bandwidth and thus a much better computational complexity. As the authors show, the method can be further optimised by recursively sampling from conditional distributions that arise from partitioning the graph via a divide-and-conquer approach; this requires the solution of sparse linear systems for which standard methods can be employed. The work [undefas] discusses the generation of Gaussian Random Fields for the simulation of noise in fluid-structure interactions on adaptive meshes for which sampling via the Fast Fourier Transformation is not applicable. They employ a variant of MGMC that is based on Fast Adaptive Composite Mesh Multigrid [undefat]. Related to this, the authors of [undefau] apply MGMC to a finite element discretisation on an unstructured mesh for a very similar physical problem.
@@
The precision operator $\mathcal{A}$ and the dimension $d$ of the problem determine the roughness of the underlying Gaussian Random Field. The commonly used operator $\mathcal{A}=(-\Delta+\kappa^{2}I)^{\alpha}$ corresponds to a class of Matern fields with $\nu=\alpha-d/2$ where $\nu$ determines the mean-square differentiability of Gaussian Random Field [undefb]. Although the discretised problem is well-defined for a finite lattice spacing, for $\nu\leq 0$ the field is so rough that the continuum limit of the probability distributions only exists in the weak sense, i.e. when evaluated against bounded linear operators. This includes the important case $\mathcal{A}=-\Delta+\kappa^{2}I$ in $d>1$ dimensions. In contrast to other methods such as the SPDE approach [undefb, undefw, undefr], this does not pose any problems for MGMC. To see this, observe that both MGMC and the SPDE approach work with the precision matrix $A$. This leads to good computational properties when $A$ is sparse, and this naturally follows if the matrix is given by a discretised differential operator of the above form. However, the SPDE approach inverts the square-root $A^{1/2}$ of the precision matrix, which corresponds to the operator $\mathcal{A}^{1/2}=(-\Delta+\kappa^{2}I)^{\alpha/2}$. If $\alpha$ is an odd integer, the SPDE approach will require the inversion of a fractional power of a sparse matrix, which is not sparse itself. This problem does not arise for MGMC which only uses $A$ and which can therefore readily be used to sample fields for $\mathcal{A}=-\Delta+\kappa^{2}I$. More generally, for arbitrary precision operators MGMC is therefore able to also cover the case where the square-root of the precision matrix is non-sparse even though the precision matrix itself is sparse. In other words, MGMC can sample from distributions with covariance operator $(-\Delta+\kappa^{2}I)^{-1}$, while additional modifications are necessary to extend the SPDE approach to this case (see e.g. [undefav]). It is also not obvious how to extend the SPDE approach to the Bayesian setting, where the precision operator is a low-rank perturbation of a differential operator. As we discuss in detail below, MGMC can be readily adapted to this case. Finally, even when the SPDE approach is applicable, it requires the inversion of a large, ill-conditioned matrix. The most efficient, versatile and parallelisable class of solvers for such matrices that also work on general domains and for non-stationary distributions are multigrid methods. However, these will require several iterations to converge, each of which will incur roughly the same cost as a single MGMC update. Thus, while a detailed comparison is beyond the scope of this paper, MGMC is in general expected to be competitive with SPDE methods.
@@
In [undefaw] a low-rank truncation of the Cholesky factorisation is used to approximately sample from Gaussian random fields. Although the method introduces an approximation, the resulting error can be controlled systematically and the method can be more efficient than the naive Cholesky approach since only a finite number of eigenmodes have to be included. However, the efficiency of the method depends on the spectrum of the covariance operator, and in particular the sufficiently rapid decay of the eigenvalues. We have found that for $\mathcal{A}=-\Delta+\kappa^{2}I$ in $d>1$ dimensions the number of required eigenmodes grows too rapidly for the method to be efficient. Again, MGMC is able to cover this case without any problems.
@@
Another advantage of the MGMC approach is that it can be readily parallelised on distributed memory machines by using well established domain decomposition techniques for the parallel implementation of multigrid solvers [undefax]. For example, the grid traversal in the Gibbs sampler that is used at every level of the grid hierarchy, needs to be performed in a red-black ordering and well-known strategies such as coarse grid aggregation [undefay] can be used on the coarser levels to take into account the fact that the number of processing units can exceed the number of unknowns. Sampling from posterior distributions in the linear Bayesian setting that we consider here, requires parallel sparse matrix vector products and scatter operations of the form $y=a(a^{\top}x)$ for vectors $a,x,y$ and again these can be implemented using well established techniques in scientific computing. Although a parallel implementation of MGMC is beyond the scope of this paper, we have every reason to believe that it will show the same excellent parallel scalability on large distributed memory machines as multigrid solvers [undefaz, undefay, undefaaa].